OpenStudy (anonymous):

An oil rig is to be built 30 miles down shore and 20 miles out to sea from a refinery. A pipe line needs to be built from the refinery to the oil rig. It costs $1000 per mile to build the pipe on land and $1500 per mile to build it at sea. The planners of this oil rig want to have the lowest cost possible for this pipe. How much would it cost to build the pipe line all the way down the shore until it is directly "below" the oil rig and then out to sea?

OpenStudy (amistre64):

are we assuming that the sea floor is flat?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

this starts out to be a minimization problem but then the question asks something different. confusing.

OpenStudy (amistre64):

10sqrt(13) * 1000 then

OpenStudy (amistre64):

the minimal distance between 2 points on a plane is a stright line

OpenStudy (amistre64):

its a pythag problem that has no real world application lol

OpenStudy (amistre64):

it seems to be asking; how much to build a pipeline along the legs of the triangle tho doesnt it

OpenStudy (anonymous):

not to me but my reading comprehension might be off.

OpenStudy (amistre64):

1 if by land; 2 if by sea..... yeah, they want the minimal cost compared to yada yada yada

OpenStudy (anonymous):

but it doesn't ask for minimum cost, even though it says "The planners of this oil rig want to have the lowest cost possible for this pipe."

OpenStudy (anonymous):

is says "How much would it cost to build the pipe line all the way down the shore until it is directly "below" the oil rig and then out to sea? 8 minutes ago"

OpenStudy (amistre64):

we have to determine at what cost the land pipe and the sea pipe are qual

OpenStudy (amistre64):

yeah, below being a pictorial term

OpenStudy (anonymous):

the shoreline is in a distance straight line from the refinery the oil rig is verticalto the shoreline and at angle between the shorline and refinery

OpenStudy (anonymous):

it looks to me as if it is just asking how much it would cost to build the pipeline 30 miles down the shore and then 20 miles out to sea. that is what the last statement says unless i am reading it incorrectly. is this for a calc class?

OpenStudy (amistre64):

25980.76 for a straight run to the oil rig

OpenStudy (amistre64):

it is; but they want some variation of that;

OpenStudy (anonymous):

30 miles down shore AND 20 miles out to sea from a refinery

OpenStudy (anonymous):

i take that to mean east 30 and then 20 north, for example. but of course i could be wrong.

OpenStudy (amistre64):

60000 if we build it at a right angle

OpenStudy (amistre64):

52426.40 if we go with a 45 at 10 feet from the start

OpenStudy (amistre64):

C(x) = s(1500) + l(1000)

OpenStudy (amistre64):

how to relate s an l? by angle perhaps?

OpenStudy (amistre64):

sin(t) = l cos(t) = 30 - cos(t) = s

OpenStudy (amistre64):

if this was speed; then the fastest speed would be?

OpenStudy (anonymous):

30mph

OpenStudy (amistre64):

along which route :)

OpenStudy (anonymous):

i dont understand your question?

OpenStudy (amistre64):

straight shot; 54083.26 10 down and 45degrees gets us: 52426.40 11 down and a distance of sqrt(19^2 + 20^2): 52379.34 12(1000) + sqrt(18^2 + 20^2)(1500) = 52360.87

OpenStudy (amistre64):

its a matter if finding the route by land and by sea that is the cheapest to build

OpenStudy (amistre64):

30(1000) + 20(1500) = 60 000 ; so just building it up the 30 then 20 is bad

OpenStudy (anonymous):

is sqrt supposed to mean log?

OpenStudy (amistre64):

log means log sqrt means square root

OpenStudy (anonymous):

what are you sqrting 11 down?

OpenStudy (amistre64):

the pipeline goes along the beach; and each foot costs so much money ($1000) in that instance, we went down 11 feet and bent the pipe to go to the rig out at sea; the sea pipe costs 1500 per foot

OpenStudy (amistre64):

ideally, when the land pipe cost matches the sea pipe cost we have the cheapest cost

OpenStudy (amistre64):

when s(1500) = l(1000) we balance out and have the cheapest cost for total pipe

OpenStudy (anonymous):

o i get it ......

OpenStudy (anonymous):

from the refinery to the oil rig would be the hypotenous of the other sides? multiplied to get the cost?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

thank you soooooo very much for your help

OpenStudy (amistre64):

can you get it from here?

OpenStudy (anonymous):

yes i can thanks

OpenStudy (anonymous):

i agree with amistre solution. however, since this is calc class you probably should (to make the teacher happy) have some sort of function to find the minimum of. so for the second part, and if you put "x" as suggested. then the distance from where the pipe enters the water to the oil rig is \[\sqrt{x^2+400}\] miles by pythagoras and the cost is $1500 per mile for a total of \[1500\sqrt{x^2+400}\] the distance from the refinery to where the pipe hits the water is 3\[30-x\] miles at a cost of $1000 per mile for a cost of \[1000(30-x)=30000-1000x\] and therefore your total cost will be \[C(x)=1500\sqrt{x^2+400}+30000-1000x\] this is the function you need to minimize

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\[C'(x)=\frac{1500x}{\sqrt{x^2+400}}-1000\] set = 0 to find the critical point get \[x=8\sqrt{5}\]

OpenStudy (anonymous):

and don't forget that the distance down the shore line is not x, but rather 30-x or about 12.11 rounded.

OpenStudy (anonymous):

i got it lol

OpenStudy (anonymous):

i did this on the fly so you should check my work. sure about the derivative though, and fairly positive about my \[x=8\sqrt{5}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

good luck.

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

lol